The ratio of male students to female students at a certain university is 5 to 4. If there is a total of 6975 students, find the number of male students and the number of female students.
Answer by mananth(16949)

male to female = 5/4
total students = 6975
..
total parts = 5+4 = 9
...
Male students = 5/9 * 6975 = 3875
female = 4/9 * 6975 = 3100
...
CHECK
3875 + 3100 = 6975